An Electronics Firm manufactures calculators at the McGregor plant and the Ennis Plant. At McGregor, the unit cost is $8.40, and the fixed cost is $7480. At Ennis, the unit cost is $7.880, and the fixed cost is $5419. The company wants the two plants to produce a combined total of 1500 calculators, and the total costs at each plant must be the same. How many should be produced at each plant?
Answer by stanbon(75887)

An Electronics Firm manufactures calculators at the McGregor plant and the Ennis Plant.
---------------------------
The company wants the two plants to produce a combined total of 1500 calculators, and the total costs at each plant must be the same. How many should be produced at each plant?
----------------------
Let number produced by McGregor be "x".
Letnumber produced by Ennis be "1500-x".
----------------------------
At McGregor, the unit cost is $8.40, and the fixed cost is $7480.
Total McGregor cost(x) = 7480 + 8.4x
--------------------------------
At Ennis, the unit cost is $7.880, and the fixed cost is $5419.
Total Ennis cost(x) = 5419 + 7.88(1500-x)
--------------------------------
EQUATION:
7480 + 8.4x = 5419 + 7.88(1500-x)
7480 + 8.4x = 5419 + 11820 - 7.88x
16.28x = 9759
x is approximately 600 (# of McGregor units if costs are going to be equal)
----------
That implies that Ennis would have to produce 900 units.
---------------
How you make the number of units equal AND the costs equal is beyond me.
=======================
Cheers,
Stan H.